for the first order reaction


A(g) ------: 2B(g) +C(g)


the initial pressure is Pa= 90 mm hg and pressure after 10 min. is 180 mm of hg. find the rate constant.ans is 1.15*10^-3


Grade:

3 Answers

vikas askiitian expert
13 years ago

A           =>           2B   +   C

p                           0         0                        (initial)                     (at t=0)

P-Po                     2Po       Po                       (final)                       (at t =10)

initial total pressure = P = 90mm

final total pressure = 180mm

 from eq adding all presures we get total pressure = P + 2Po

P +  2Po = 180

90 + 2Po = 180

   Po = 45

for A , initially pressure  = P= 90

        final pressure = P-Po =  90-45 = 45

rate constant (k) = 2.303 log(Pi/Pf) / t

                        = 2.303 log90/45 / t

                        =(2.303)log2/10

                        k = 0.069 min-1

this value of rate constant is in min-1 , 1min = 60sec so

                   K = 0.069/60 sec-1  = 1.15*10-3 sec-1

approve if u like my ans

SAGAR SINGH - IIT DELHI
13 years ago

Dear student.,

Write the rate of reation in terms of pressure

r=(2.303log P/Po)/time

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Sagar Singh

B.Tech, IIT Delhi

Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

A => 2B + C
p 0 0 (initial) (at t=0)
P-Po 2Po Po (final) (at t =10)
initial total pressure = P = 90mm
final total pressure = 180mm
from eq adding all presures we get total pressure = P + 2Po
P + 2Po = 180
90 + 2Po = 180
Po = 45
for A , initially pressure = P= 90
final pressure = P-Po = 90-45 = 45
rate constant (k) = 2.303 log(Pi/Pf) / t
= 2.303 log90/45 / t
=(2.303)log2/10
k = 0.069 min-1
this value of rate constant is in min-1 , 1min = 60sec so
K = 0.069/60 sec-1 = 1.15*10-3 sec-1

Thanks and Regards